# Concat-ID: Towards Universal Identity-Preserving Video Synthesis

### v2

@@ Abstract
We present Concat-ID, a unified framework for identity-preserving video generation. Concat-ID employs Variational Autoencoders to extract image features, which are concatenated with video latents along the sequence dimension, leveraging solely 3D self-attention mechanisms without the need for additional modules. A novel cross-video pairing strategy and a multi-stage training regimen are introduced to balance identity consistency and facial editability while enhancing video naturalness. Extensive experiments demonstrate Concat-ID’s superiority over existing methods in both single and multi-identity generation, as well as its seamless scalability to multi-subject scenarios, including virtual try-on and background-controllable generation. Concat-ID establishes a new benchmark for identity-preserving video synthesis, providing a versatile and scalable solution for a wide range of applications.
@@ 1 Introduction
Identity-preserving video generation, which seeks to create human-centric videos of a specific identity accurately matching a user-provided face image, has recently gained significant attention, as evidenced by the success of commercial tools such as Vidu [23] and Pika [19]. A primary challenge in this field is achieving a balance between maintaining identity consistency and enabling facial editability. Prior work [28, 9, 30, 13] fails to effectively preserve identity despite utilizing special face encoders and incorporating extra adapters to mitigate cross-modal disparities.
To mitigate this limitation, some approaches [29, 4] substitute the spatially aligned reference image in pre-trained image-to-video models [2, 27] with facial images, leading to a significant improvement in identity consistency. However, they still face challenges in preventing the replication of facial expressions from the reference image. Moreover, the supplementary modules and parameters introduced by these methods contribute to increased complexity in both model training and inference. Furthermore, academic research on identity-preserving video generation that incorporates multiple identities and subjects remains scarce, leaving it notably behind commercial applications [19, 23].
In this work, we introduce Concat-ID, a concise, effective, and versatile framework for identity-preserving video generation. By unifying the model architecture, data processing, and training procedure, Concat-ID not only achieves single-identity video generation but also seamlessly integrates multiple identities and accommodates diverse subjects. Specifically, Concat-ID employs Variational Autoencoders (VAEs) to extract image features, which are then concatenated with video latents along the sequence dimension. This approach leverages solely 3D self-attention mechanisms, which are inherently present in state-of-the-art video generation models, thereby eliminating the need for extra modules and parameters. Furthermore, to effectively balance identity consistency and facial editability while enhancing video naturalness, we develop a novel cross-video pairing strategy and a multi-stage training regimen.
@@
  • Concat-ID uses VAEs to extract image features, which are concatenated with video latents along the sequence dimension, relying solely on 3D self-attention mechanisms.
@@ 3 Preliminary
Figure 1: The architecture of Concat-ID. We utilize a VAE to extract image latents from reference images and concatenate them at the end of the video latents along the sequence dimension. Concat-ID relies solely on 3D self-attention, which are commonly present in state-of-the-art video generation models, without introducing additional modules and parameters.
@@ 4 Concat-ID
Given a reference image containing a human face, our goal is to generate identity-preserving videos based on user-provided text prompts, while also enabling the integration of additional identities or subjects. To address this challenge, we propose Concat-ID, a concise, effective, and versatile framework. As illustrated in Fig. 1, we introduce a unified architecture for extracting and injecting features from any number of identities and subjects without requiring extra modules and parameters (see Sec. 4.1). To balance identity consistency and facial editability while enhancing video naturalness, we further construct cross-video pairs as training data (see Sec. 4.2) and propose a novel multi-stage training strategy (see Sec. 4.3).
@@ 4.1 A unified architecture
We focus on designing a unified model architecture for identity feature fusion that is compatible with state-of-the-art text-to-video and image-to-video models, and readily extended to multi-identity and multi-subject scenarios. Revisiting the role of VAEs, we recognize their ability to compress conditioning images into the same latent space as $\mathbf{Z}$. Consequently, unlike prior works [28, 9, 30, 13] that introduce extra modules to mitigate distribution differences between image features and $\mathbf{Z}$, our denoising transformer $\bm{\epsilon}_{\bm{\theta}}$ can inherently interpret these features due to pre-training on related videos and images. Based on this insight, we adopt the VAE as our feature extractor.
@@
where $\text{Concat}(\cdot,\cdot,\cdots)$ denotes concatenation along the sequence dimension and $\mathbf{Z^{\prime}}\in\mathbb{R}^{(N+M)\times HW\times C}$. As shown in Fig. 1, this feature injection through concatenation is compatible with any video generation model that utilizes 3D self-attention, which are generally present in state-of-the-art video generation models. Since $\mathbf{Z}$ and $\mathbf{c}_{i}$ are in the same latent space, $\bm{\epsilon}_{\bm{\theta}}$ can seamlessly integrate identity-preserving features without the need for additional modules and parameters to address cross-modal disparities.
Concatenating $\mathbf{Z}$ and $\mathbf{c}_{i}$ along the channel dimension is another direct method for feature injection, as employed in ConsisID [29] and Ingredients [4]. However, this strategy introduces artifacts (see Fig. 3 and Fig. 4) due to spatial misalignment between face images and video latents. In contrast, by leveraging a 3D self-attention mechanism, our sequence concatenation promotes spatial interactions without compromising the quality of any generated frame. Furthermore, it scales efficiently to handle multi-identity and multi-subject scenarios (see LABEL:fig:examples). For instance, when $M=2$, where $\mathbf{I}_{1}$ is a face image and $\mathbf{I}_{2}$ represents clothing, our method can extract clothing features $\mathbf{c}_{2}$ using $\mathcal{E}$ and achieve virtual try-on while preserving the given identity $\mathbf{I}_{1}$.
@@ 4.2 Data construction
Personalized image generation can also synthesize reference images with the same identity as given videos but varied facial expressions, as demonstrated in [10, 20]. However, this approach incurs high computational costs, particularly for large-scale image-video pairs. Additionally, existing personalized generation methods [28, 7, 33] often struggle to preserve detailed facial features, which limits their effectiveness. In contrast, as shown in Fig. 2(b), our retrieval-based method efficiently gathers a large-scale set of real reference images that accurately match the identity of corresponding videos while exhibiting diversity across multiple dimensions, such as facial expressions, hairstyles, lighting conditions, and other identity-irrelevant factors.
@@
Notably, trade-off pairs are not strictly necessary and we introduce them to achieve an optimal balance between identity consistency and facial editability. As shown in Fig. 2(c), they exhibit greater similarity than cross-video pairs, which can enhance identity consistency in trained models. Moreover, thresholds for cross-video pairing are set empirically.
@@ 4.3 Training strategy
Building on our innovative data construction, we introduce a multi-stage training process: pre-training stage, cross-video fine-tuning, and trade-off fine-tuning. In the pre-training stage, we optimize a text-to-video model on $\mathcal{S}_{\textrm{pre}}$ to map facial details into generated videos. This self-supervised training method may constrain certain generated video frames to adhere strictly to the given condition images, potentially degrading the editability of facial expressions and the overall naturalness. The cross-video fine-tuning on $\mathcal{S}_{\textrm{cross}}$, using image-video pairs derived from different videos, can alleviate this issue. However, we observe that this fine-tuning enhances facial editability at the expense of identity fidelity (see Sec. 5.4). To balance fidelity and editability while ensuring high-degree motion and high artistic quality, we ultimately fine-tune the model on $\mathcal{S}_{\textrm{trade}}$.
Throughout all training stages, we proportionally scale, pad, and center-crop images to match the video resolution. To ensure the model focuses on facial regions during training and prevents background leakage during inference, we segment and drop the background of reference images [26]. Additionally, to improve robustness and generalization, we introduce random noise to reference images during training, while omitting this noise during inference. To further differentiate the image latent $\mathbf{c}_{i}$ from the video latent $\mathbf{Z}$ and distinguish between different $\mathbf{c}_{i}$, we extend 3D-RoPE to incorporate multiple reference images along the sequence dimension. Owing to the simplicity and efficiency of Concat-ID in both data construction and model architecture, our training strategy can seamlessly scale to multi-identity and multi-subject scenarios. Moreover, we establish that single-identity pre-training facilitates enhanced identity preservation in these downstream tasks (see Tab. 2).
@@ 5.1 Experimental settings
Baselines. For a comprehensive comparison, we use three representative open-source approaches as baselines. (1) Single-Identity personalization methods: ID-Animator [9] and ConsisID [29]. (2) Multi-Identity personalization methods: Ingredients [4]. ID-Animator, ConsisID, and Ingredients all incorporate additional adapters and auxiliary loss functions to enhance identity consistency. Notably, to the best of our knowledge, Ingredients [4] is the only available open-source method for multiple identities, and no open-source methods currently exist for multiple subjects.
@@ 5.2 Main results
Table 1: Quantitative results for single-identity and multi-identity generation. $\ddagger$ indicates corresponding methods introduce additional adapters and auxiliary loss. Concat-ID achieves superior identity consistency and facial editability while maintaining better or comparable text alignment relative to the baselines.
@@ 5.3 Multiple identities and subjects
The architecture, data construction, and training strategy of Concat-ID allow for its seamless extension to multi-identity and multi-subject scenarios, sustaining good performance.
@@ 6 Conclusions
In this paper, we introduce Concat-ID, a unified framework for identity-preserving video generation. Concat-ID relies solely on 3D self-attention mechanisms, which are commonly used in state-of-the-art video generation models, without introducing additional modules. We also present a novel cross-video pairing strategy and a multi-stage training regimen to balance identity consistency and facial editability while enhancing video naturalness. Thanks to its architecture, data construction, and training strategy, Concat-ID achieves effective video synthesis across single-identity, multi-identity, and multi-subject scenarios.
Limitations. Due to limitations in the base model’s capabilities, we do not compare our method with closed-source commercial tools. Currently, we utilize VAEs solely as feature extractors, relying on the model’s inherent ability to process low-level features. Similar to common video generation models, our approach faces challenges in preserving the integrity of human body structures, such as the number of fingers, when handling particularly complex motions.
@@ A.2 Implementation details
In the first stage, we randomly select one reference image from a set of five for each video. Traditional data augmentation techniques, such as flipping, are not used for face images, as they can cause data augmentation leakage [16], leading the model to learn the augmented data distribution rather than the original distribution. For instance, horizontal flipping may result in incorrectly mirrored faces in generated videos. For inference, by default, we set the classifier-free guidance scale to 6, 5, and 5 for the first, second, and third stages, respectively.
@@ B.2 Multi-subject scenarios
To maximize model performance, we independently train different specialized models for specific tasks. Developing a comprehensive model capable of addressing multiple tasks simultaneously is left for future work.

### v3

@@ Abstract
We present Concat-ID, a unified framework for identity-preserving video generation. Concat-ID employs variational autoencoders to extract image features, which are then concatenated with video latents along the sequence dimension. It relies exclusively on inherent 3D self-attention mechanisms to incorporate them, eliminating the need for additional parameters or modules. A novel cross-video pairing strategy and a multi-stage training regimen are introduced to balance identity consistency and facial editability while enhancing video naturalness. Extensive experiments demonstrate Concat-ID’s superiority over existing methods in both single and multi-identity generation, as well as its seamless scalability to multi-subject scenarios, including virtual try-on and background-controllable generation. Concat-ID establishes a new benchmark for identity-preserving video synthesis, providing a versatile and scalable solution for a wide range of applications.
@@ 1 Introduction
Identity-preserving video generation, which seeks to create human-centric videos of a specific identity accurately matching a user-provided face image, has recently gained significant attention, as evidenced by the success of commercial tools such as Vidu [23] and Pika [19].
A primary challenge in this field is achieving a balance between maintaining identity consistency and enabling facial editability. Prior work [28, 9, 30, 13] fails to effectively preserve identity despite utilizing special face encoders and incorporating extra adapters to mitigate cross-modal disparities. To mitigate this limitation, some approaches [29, 4] substitute the spatially aligned reference image in pre-trained image-to-video models [2, 27] with facial images, leading to a significant improvement in identity consistency. However, they still face challenges in preventing the replication of facial expressions from the reference image. Moreover, the supplementary modules and parameters introduced by these methods contribute to increased complexity in both model training and inference.
In this work, we introduce Concat-ID, a concise, effective, and versatile framework for identity-preserving video generation. By unifying the model architecture, data processing, and training procedure, Concat-ID not only achieves single-identity video generation but also seamlessly integrates multiple identities and accommodates diverse subjects. Specifically, Concat-ID employs Variational Autoencoders (VAEs) to extract image features, which are then concatenated with video latents along the sequence dimension. This approach relies exclusively on 3D self-attention mechanisms, which are inherently present in state-of-the-art video generation models, to incorporate image features, thereby eliminating the need for extra modules or parameters. Furthermore, to effectively balance identity consistency and facial editability while enhancing video naturalness, we develop a novel cross-video pairing strategy and a multi-stage training regimen.
@@
  • Concat-ID utilizes VAEs to extract image features and integrates them via inherent 3D self-attention mechanisms, without introducing additional parameters or modules.
@@ 3 Preliminary
Figure 1: The architecture of Concat-ID. We utilize VAEs to extract image latents from reference images and concatenate them at the end of the video latents along the sequence dimension. Concat-ID relies solely on 3D self-attention mechanisms, which are commonly present in state-of-the-art video generation models, to integrate image features without adding extra modules or parameters.
@@ 4 Concat-ID
Given a reference image containing a human face, our goal is to generate identity-preserving videos based on user-provided text prompts, while also enabling the integration of additional identities or subjects. To address this challenge, we propose Concat-ID, a concise, effective, and versatile framework. As illustrated in Fig. 1, we introduce a unified architecture for extracting and injecting features from any number of identities and subjects without requiring extra modules or parameters (see Sec. 4.1). To balance identity consistency and facial editability while enhancing video naturalness, we further construct cross-video pairs as training data (see Sec. 4.2) and propose a novel multi-stage training strategy (see Sec. 4.3).
@@ 4.1 A unified architecture
We focus on designing a unified model architecture capable of extracting and fusing the identity feature and readily extendable to multi-identity and multi-subject scenarios. Revisiting the role of VAEs, we recognize their ability to compress conditioning images into the same latent space as the video latent $\mathbf{Z}$. Consequently, our denoising transformer $\bm{\epsilon}_{\bm{\theta}}$ can inherently interpret these features. Based on this insight, we adopt the VAE as our feature extractor.
@@
where $\text{Concat}(\cdot,\cdot,\cdots)$ denotes concatenation along the sequence dimension and $\mathbf{Z^{\prime}}\in\mathbb{R}^{(N+M)\times HW\times C}$. As shown in Fig. 1, this feature injection through concatenation is compatible with any video generation model that utilizes 3D self-attention, which are generally present in state-of-the-art video generation models. Since $\mathbf{Z}$ and $\mathbf{c}_{i}$ are in the same latent space, $\bm{\epsilon}_{\bm{\theta}}$ can seamlessly integrate identity-preserving features without the need for additional modules or parameters to address cross-modal disparities.
Concatenating $\mathbf{Z}$ and $\mathbf{c}_{i}$ along the channel dimension is another direct method for feature injection, as employed in ConsisID [29] and Ingredients [4]. However, this strategy introduces artifacts (see Fig. 3 and Fig. 4) due to spatial misalignment between face images and video latents. In contrast, by leveraging a 3D self-attention mechanism, our sequence concatenation promotes spatial interactions without compromising the quality of any generated frame. Furthermore, it scales efficiently to handle multi-identity and multi-subject scenarios (see LABEL:fig:examples).
@@ 4.2 Data construction
Personalized image generation can also synthesize reference images with the same identity as given videos but varied identity-irrelevant factors, as demonstrated in [10, 20]. However, this approach incurs high computational costs, particularly for large-scale image-video pairs. Additionally, existing personalized generation methods [28, 7, 33] often struggle to preserve detailed facial features, which limits their effectiveness. In contrast, as shown in Fig. 2(b), our retrieval-based method efficiently gathers a large-scale set of real reference images that accurately match the identity of corresponding videos while exhibiting diversity across multiple dimensions, such as facial expressions, hairstyles, lighting conditions, and other identity-irrelevant factors.
@@ 4.3 Training strategy
Building on our innovative data construction, we introduce a multi-stage training process: pre-training stage, cross-video fine-tuning, and trade-off fine-tuning. In the pre-training stage, we optimize a text-to-video model on $\mathcal{S}_{\textrm{pre}}$ to map facial details into generated videos. This self-supervised training method may constrain certain generated video frames to adhere strictly to the given condition images, potentially degrading the editability of facial expressions and the overall naturalness. The cross-video fine-tuning on $\mathcal{S}_{\textrm{cross}}$, using image-video pairs derived from different videos, can alleviate this issue. However, we observe that this fine-tuning enhances facial editability at the expense of identity fidelity (see Sec. 5.4).
A simple strategy to further balance fidelity and editability is to mix pre-trained pairs and cross-video pairs in a 1:1 ratio, a similar method adopted by Movie-Gen [20]. However, our initial experiments suggest that this approach results in unstable training due to varying identity consistency between pre-trained pairs and cross-video pairs. To address this issue while ensuring high-degree motion and high artistic quality, we ultimately fine-tune the model on $\mathcal{S}_{\textrm{trade}}$.
Throughout all training stages, we proportionally scale, pad, and center-crop images to match the video resolution. To ensure the model focuses on facial regions during training and prevents background leakage during inference, we segment and drop the background of reference images [26]. Additionally, to improve robustness and generalization, we introduce random noise to reference images during training, while omitting this noise during inference. To further differentiate the image latent $\mathbf{c}_{i}$ from the video latent $\mathbf{Z}$ and distinguish between different $\mathbf{c}_{i}$, we extend 3D-RoPE to incorporate multiple reference images along the sequence dimension. Specifically, we introduce a temporal bias $N$ to define the 3D position of a token $\mathbf{t}_{h,w}$ in $\mathbf{c}_{i}$:
[…]
Owing to the simplicity and efficiency of Concat-ID in both data construction and model architecture, our training strategy can seamlessly scale to multi-identity and multi-subject scenarios. Moreover, we establish that single-identity pre-training facilitates enhanced identity preservation in these downstream tasks (see Tab. 2).
@@ 5.1 Experimental settings
Baselines. For a comprehensive comparison, we use three representative open-source approaches as baselines. (1) Single-identity personalization methods: ID-Animator [9] and ConsisID [29]. (2) Multi-identity personalization methods: Ingredients [4]. ID-Animator, ConsisID, and Ingredients all incorporate additional adapters and auxiliary loss functions to enhance identity consistency. Notably, Concat-ID, ConsisID, and Ingredients are all built upon the same video model, CogVideoX-5B.
@@ 5.2 Main results
[…]
@@ 5.3 Multiple identities and subjects
We demonstrate that the architecture, data construction, and training strategy of Concat-ID make it seamlessly extendable to multi-identity and multi-subject scenarios.
@@ 6 Conclusions
In this paper, we introduce Concat-ID, a unified framework for identity-preserving video generation. Concat-ID relies solely on 3D self-attention mechanisms, which are commonly used in state-of-the-art video generation models, without introducing additional modules or parameters. We also present a novel cross-video pairing strategy and a multi-stage training regimen to balance identity consistency and facial editability while enhancing video naturalness. Thanks to its architecture, data construction, and training strategy, Concat-ID can scale seamlessly to multi-identity and multi-subject scenarios.
Limitations. Similar to common video generation models, our approach faces challenges in preserving the integrity of human body structures, such as the number of fingers, when handling particularly complex motions. In this paper, we focus on the single-identity scenario, and further improvement and evaluation of Concat-ID’s performance in multiple-identity and multi-subject scenarios is left for future work.
@@ A.2 Implementation details
In the first stage, we randomly select one reference image from a set of five for each video. Traditional data augmentation techniques, such as flipping, are not used for face images, as they can cause data augmentation leakage [16], leading the model to learn the augmented data distribution rather than the original distribution. For instance, horizontal flipping may result in incorrectly mirrored faces in generated videos.
@@ A.3 Baselines
A.4 Training cost
The first, second, and third stages of training in the single-identity scenario required 3,260, 2,104, and 135 NVIDIA H800 GPU hours, respectively, with the cost of the third stage being negligible.
@@ B.2 Multi-subject scenarios
In this paper, we focus on the single-identity scenario, and improving the performance of Concat-ID in multiple-identity and multi-subject settings is left for future work. To maximize model performance, we independently train different specialized models for specific tasks. The development of a comprehensive model capable of addressing multiple tasks simultaneously remains a direction for future research.
@@ Appendix C Ablation study
Trade-off pairs can naturally enhance the identity consistency of Concat-ID, as they maintain better alignment between reference images and videos compared to cross-video pairs. An interleaved training strategy—alternating between Stage I for improving identity and Stage II for enhancing editability—can also achieve a favorable trade-off, a method similarly adopted in Imagine-yourself [10]. However, our multi-stage training approach achieves an optimal balance just by adding a third stage where we carefully control identity consistency and sample quantity, showing that a simple design can be highly effective.